# Hybrid inflation and gravitational waves from accidentally light scalars

### v1

@@ Abstract
We construct a hybrid-inflation model where the inflaton potential is generated radiatively, as gauge symmetries guarantee it to be accidentally flat at tree level. The model can be regarded as a small-field version of Natural Inflation, with inflation ending when the mass of a second scalar, the waterfall field, turns tachyonic. This provides a minimal, robust realisation of hybrid inflation, which predicts specific correlations among CMB observables. Tachyonic preheating leads to the production of gravitational waves which, for a low inflationary scale, might be detected by upcoming experiments. Simple variations of the model can give rise to topological defects, such as unstable domain walls. Their dynamics produces a stochastic gravitational-wave background, which can be compatible with the recent detection by pulsar timing arrays.
@@ 1 Introduction
Field-theoretical models of cosmic inflation can be roughly divided into two classes. In large-field models, the field excursion of the inflaton is transplanckian, while in small-field models it remains below the Planck scale. When allowing for large field excursions, very simple shapes of the potential can give rise to inflation (although many are now excluded by increasingly accurate CMB measurements [1, 2]). However, it is generally dubious if large-field models can be trustable effective field theories, without being embedded in a complete quantum gravity framework. Small-field models, on the other hand, generally require more involved field-theoretic constructions. In particular, some mechanism should be included to protect the inflaton potential from radiative corrections. Once this is achieved, effects from transplanckian physics can be argued to be under control in the effective field theory. The only necessary fine-tuning left is that of the cosmological constant after inflation, and potentially of the initial conditions leading to an inflationary phase.
@@
In small-field models, the flatness of the inflaton potential could be ensured, for example, by (spontaneously broken) supersymmetry, or by an (approximate) shift symmetry if the inflaton is a pseudo Nambu-Goldstone boson (pNGB). In the present work, we propose to employ a novel mechanism to this end: the inflaton is an accidentally massless scalar field, i.e. a scalar whose potential is constrained by symmetry to be flat up to loop corrections and higher-dimensional operators [3]. Similarly to the pNGB case, a small slope is generated radiatively.
@@
At first sight, our inflaton potential is very similar to that of Natural Inflation (NI), where the inflaton is realized as a $\mathrm{U}(1)$ pNGB [4]. Minimal NI, however, is a large-field model, and moreover is excluded by Planck data [2]. Small-field NI needs additional fields responsible for ending inflation, and such models of “hybrid NI” turn out to be difficult to construct, since the couplings between the additional fields and the inflaton should not spoil the flatness of the potential (although this can be achieved with suitable discrete symmetries [5, 6]).
@@
Our model, by contrast, is simple and economical; in the minimal case the symmetry is simply a gauged $\mathrm{SO}(3)$ times a $\mathbb{Z}_{2}$ parity. Indeed, models with accidentally light scalars [3] lend themselves particularly well to hybrid inflation model-building. This is because they feature an almost-flat inflaton direction by construction, yet the other scalar fields couple to the inflaton at the tree level, which allows to naturally end inflation by a second field becoming tachyonic.
@@
We will compute the Cosmic Microwave Background (CMB) observables as a function of the model parameters, and study their correlations. We will also present naturalness arguments to identify the preferred region of parameters, and derive the corresponding predictions.
@@
While naturalness would prefer an inflationary scale not too far from the Planck scale, lower-scale scenarios are more interesting for the potential observation of gravitational wave (GW) signals. In hybrid models such as ours, inflation ends because a steep tachyonic “waterfall” direction appears in field space [7], which leads to a transition from the inflationary slow-roll regime to a phase of tachyonic preheating [8, 9]. In this phase, GWs can be produced, and may be within the reach of future detectors, if the inflationary scale is low. In addition, spontaneous symmetry breaking (SSB) in the inflaton-waterfall potential may lead to the production of topological defects, which also source GWs. We will present slightly extended versions of the minimal model, which produce cosmic strings or domain walls after the end of inflation. These scenarios are phenomenologically interesting only if the inflationary scale is sufficiently low. Then, the annihilations of topological defects produce GWs which can be within experimental reach.
@@
We will compute these various GW signals, study their correlations with the CMB observables, and compare them with present constraints from LIGO-Virgo-KAGRA [10] and NANOGrav [11], as well as with the expected sensitivity of future GW detectors.
@@
Both large-field and small-field models are the subject of several swampland conjectures, which aim to characterize the set of effective field theories that are supposedly incompatible with general principles of quantum gravity; see [12] for a review. Specifically, models with transplanckian field excursions are concerned by the Distance Conjecture [13]. This conjecture states roughly that, at large distances in field space, an infinite tower of fields becomes light, invalidating the effective field theory. On the other hand, slow-roll potentials with subplanckian field excursions and positive vacuum energy are claimed to be ruled out by various (arguably more speculative) de Sitter conjectures [14, 15, 16]. While these conjectures are certainly interesting, none of them is firmly established. For this work we therefore assume that single-field slow-roll inflation is not per se in the swampland, and that small-field models, in particular, can be viable effective field theories of inflation.
@@ 2 Hybrid inflation with an accidentally light scalar
Here $T^{a}_{BC}$ are the $\mathrm{SO}(3)$ generators in the ${\bf 5}$-plet representation ($a=1,2,3$ and $A,B,C=1\ldots 5$). The continuous symmetry of this potential is strictly $G=\mathrm{SO}(3)$: there is no larger custodial-type symmetry present. More precisely, the symmetry of the kinetic term is $\mathrm{SO}(8)$, which is explicitly broken to $\mathrm{SO}(5)_{\phi}\times\mathrm{SO}(3)_{\chi}$ by the mass terms and the $\lambda_{\phi,\chi}$ and $\epsilon$ quartics. The $\zeta$ coupling then breaks the symmetry explicitly to $\mathrm{SO}(3)$.
@@
At a special point along the flat direction, corresponding to a null eigenvector of one of the $G$ generators, an O$(2)=\mathrm{SO}(2)\times\mathbb{Z}_{2}$ subgroup of $G$ is preserved. At this point there are only two NGBs, while not one but two scalars remain accidentally massless at the tree level; the second one is the scalar which is absorbed by the $\mathrm{SO}(2)$ gauge boson outside the special point. This point turns out to be a minimum of the one-loop effective potential.
For concreteness, consider an explicit parametrization of the tree-level flat direction by writing the isospin-2 field $\phi$ as a $3\times 3$ traceless symmetric matrix $\Phi=\phi_{A}\lambda_{A}$. Here $\lambda_{A}$ are the the symmetric Gell-Mann matrices $\lambda_{1,3,4,6,8}$. By an $\mathrm{SO}(3)$ transformation, $\Phi$ can be diagonalized, such that its VEV becomes
@@
$m_{\chi^{1,2}}^{2}=-\tilde{\mu}_{\chi}^{2}+2\,v^{2}\zeta\,\sin^{2}\left(\alpha\pm\frac{\pi}{3}\right)\,,\quad m_{\chi^{3}}^{2}=-\tilde{\mu}_{\chi}^{2}+2\,\zeta v^{2}\sin^{2}\alpha$ (4)
@@
When including one-loop corrections to the potential, the vacuum degeneracy is lifted, the physical vacuum is at the point $\alpha=0$, and the two tree-level-massless scalars become massive. The only massless state is the gauge boson of the unbroken $\mathrm{SO}(2)$. However, there remains a one-loop mass hierarchy between the six heavy scalar modes having obtained their masses at the tree level, and the two accidentally light ones with one-loop masses. We emphasize that the latter are calculable and finite as functions of $\mu_{\phi}^{2}$ and $\mu_{\chi}^{2}$, and therefore they are quadratically sensitive to quantum corrections at the cutoff scale only at two loops (since contributions to the mass parameters themselves are quadratically divergent at one loop), much as in Little Higgs models (for a review of the latter see e.g. Ref. [17]).
Focusing on the one-loop effective potential along the tree-level flat direction, one finds that it is given by  [^1] To calculate the effective potential, we are using the renormalization prescription detailed in Ref. [3]. As a periodic and even function, $V_{\rm eff}(a)$ can be expanded into its Fourier cosine series $\sum_{n}c_{n}\cos(na/f)$.
$V_{\rm eff}(a)=V_{0}-M^{4}\cos\frac{a}{f}+(\text{higher harmonics})~{},$ (6)
@@
$M^{4}=\frac{1}{160\,\pi^{2}}\left[9\,g^{4}\,v^{4}+2\,\frac{\tilde{\mu}_{\chi}^{10}}{\zeta^{3}\,v^{6}}\left(F\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)-T_{F,5}\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)\right)\right]~{},$ (7)
@@
The shape of the one-loop potential, therefore, closely resembles what one would obtain if $a$ were the pNGB of an approximate global $\mathrm{U}(1)$ symmetry. Identifying $a$ with the inflaton and setting $V_{0}=M^{4}$, one obtains a realization of the famous Natural Inflation (NI) model [4]. However, vanilla NI is excluded by Planck data [2]. Simple modifications of the minimal NI model, such as a non-minimal coupling to gravity [18, 19, 20] or allowing for some exotic phase of reheating [21] can improve the fit, but in the preferred parameter region one finds $f\gtrsim M_{\mathrm{P}}$. [^3] We designate by $M_{\mathrm{P}}$ the reduced Planck mass, $M_{\mathrm{P}}\equiv(8\pi G_{\mathrm{\scriptscriptstyle{N}}})^{-1/2}=2.4\times 10^{18}\,{\rm GeV}$. As already discussed, transplanckian field excursions, apart from being constrained by the Swampland Distance Conjecture, imply that the effects of higher-dimensional operators are not under control (although models have been constructed which address this issue [22, 23]). To overcome these problems, it is desirable to identify a variant of the model where the field excursion is sub-Planckian, i.e. a small-field model of inflation.
In a small-field version of NI, the inflationary potential should be dominated by the constant term $V_{0}$ in the vacuum energy. This implies that slow-roll is possible for $f<M_{\mathrm{P}}$, but it also implies that slow-roll never ends on its own because the slow-roll parameters never grow to become ${\cal O}(1)$, and that the vacuum energy density $V_{0}$ somehow needs to be converted into radiation eventually. To this end, one may extend the model into one of hybrid inflation [7]: add a second scalar field, the waterfall field, which is heavy during inflation but becomes tachyonic for some value of the inflaton, triggering the end of inflation and the start of preheating.
@@
So far the role of the field $\chi$ was solely to provide a source for loop corrections to the effective potential which lift the flatness of the inflaton potential. We will now identify $\chi$ with the waterfall field by choosing the model parameters such that the true vacuum of the model is no longer at $\langle\chi\rangle=0$ and $\langle\phi\rangle\neq 0$ but rather at $\langle\phi\rangle=0$ and $\langle\chi\rangle\neq 0$. Our previous discussion of the vacuum structure applies as long as the effective $\chi$ masses, given in Eq. (4), are all positive, regardless of the sign of the mass parameter $\tilde{\mu}_{\chi}^{2}$ of Eq. (5). Even for $\tilde{\mu}_{\chi}^{2}>0$, the effective potential is still given by Eqs. (6) and (7) upon replacing $M^{4}$ by its real part. It is only when a tree-level mass becomes tachyonic that our computation of $V_{\rm eff}$ can no longer be trusted.
@@
In this regime, the inflaton $a$ starts rolling close to $\alpha=\pi/6$ or equivalently $a=\pi f$, where all scalars except $a$ have positive tree-level masses. In particular, $\chi$ is stabilized at $\chi=0$. For $\alpha>\alpha_{c}$ the one-loop effective potential Eq. (6) can be trusted, and provides a small slope which the inflaton rolls down. As the inflaton crosses the point $\alpha_{c}$, the tree-level vacuum will no longer be given by $\chi=0$ and Eq. (2), since a tree-level tachyon develops in the $\chi^{3}$ direction. Slow-roll ends as the tachyonic $\chi$ mass grows large, and $\chi$ modes are being excited. Eventually, the potential energy of the inflaton is converted into energy in the $\chi$ sector. In Fig. 1 we provide a sketch of the potential along the $a$ (inflaton) and $\chi^{3}$ (waterfall) directions, for the sake of illustration.
Figure 1: Pictorial representation of the inflationary potential, normalised to the inflationary scale $V_{0}$. Close to $a=\pi f$, where inflation begins, the potential along the $\chi^{3}$ direction (purple lines) has positive and large curvature, stabilizing the inflationary trajectory (red line). At the $a=a_{*}$ the pivot scale crosses the Hubble horizon, $RH=k_{*}$ (with $R$ the scale factor). As the inflaton slow-rolls down its potential, $m_{\chi^{3}}^{2}$ decreases until it becomes 00 at the critical point $a_{c}$, where inflation ends. For $a<a_{c}$, the effective potential cannot be trusted, as the accidentally flat direction is no longer a minimum of the tree-level potential. A value of $a_{*}/f$ ($a_{c}/f$) in the green (blue) region would correspond to a tuning $\gtrsim 10\%$ on the initial (final) point of observable inflation, and therefore it is theoretically disfavoured.
Reheating should then proceed via some coupling, such as a Higgs portal coupling, between $\chi$ and the Standard Model; in this work we will not explore reheating in detail. It is easily checked that, for suitable values of the couplings, the true vacuum of the model is located at
@@
which is the (tree-level) potential energy difference between the inflationary field configuration and the true vacuum. In the following we will assume that the second term in Eq. (11) is not only subdominant but even negligible.

### v2

@@ Abstract
We construct a hybrid-inflation model where the inflaton potential is generated radiatively, as gauge symmetries guarantee it to be accidentally flat at tree level. The model can be regarded as a small-field version of Natural Inflation, with inflation ending when the mass of a second scalar, the waterfall field, turns tachyonic. This provides a minimal, robust realization of hybrid inflation, which predicts specific correlations among Cosmic Microwave Background observables. Tachyonic preheating leads to the production of gravitational waves which, for a low inflationary scale, might be detected by upcoming experiments. Simple variations of the model can give rise to topological defects, such as unstable domain walls. Their dynamics produces a stochastic gravitational-wave background, which can be compatible with the recent detection by pulsar timing arrays.
@@ 1 Introduction
Field-theoretical models of cosmic inflation can be roughly divided into two classes. In large-field models, the field excursion of the inflaton is transplanckian, while in small-field models it remains below the Planck scale. When allowing for large field excursions, very simple shapes of the potential can give rise to inflation (although many are now excluded by increasingly accurate Cosmic Microwave Background (CMB) measurements [1, 2]). However, it is generally dubious if large-field models can be trustable effective field theories, without being embedded in a complete quantum gravity framework. Small-field models, on the other hand, generally require more involved field-theoretic constructions. In particular, some mechanism should be included to protect the inflaton potential from radiative corrections. Once this is achieved, effects from transplanckian physics can be argued to be under control in the effective field theory. The only necessary fine-tuning left is that of the cosmological constant after inflation, and potentially of the initial conditions leading to an inflationary phase.
@@
In small-field models, the flatness of the inflaton potential could be ensured, for example, by (spontaneously broken) supersymmetry, or by an (approximate) shift symmetry if the inflaton is a pseudo Nambu-Goldstone boson (pNGB). In the present work, we propose to employ a novel mechanism to this end: The inflaton is an accidentally massless scalar field, i.e., a scalar whose potential is constrained by symmetry to be flat up to loop corrections and higher-dimensional operators [3]. Similarly to the pNGB case, a small slope is generated radiatively.
@@
At first sight, our inflaton potential is very similar to that of Natural Inflation (NI), where the inflaton is realized as a $\mathrm{U}(1)$ pNGB [4]. Minimal NI, however, is a large-field model, and, moreover, is excluded by Planck data [2]. Small-field NI needs additional fields responsible for ending inflation, and such models of “hybrid NI” turn out to be difficult to construct, since the couplings between the additional fields and the inflaton should not spoil the flatness of the potential (although this can be achieved with suitable discrete symmetries [5, 6]).
@@
Our model, by contrast, is simple and economical; in the minimal case, the symmetry is simply a gauged $\mathrm{SO}(3)$ times a $\mathbb{Z}_{2}$ parity. Indeed, models with accidentally light scalars [3] lend themselves particularly well to hybrid inflation model-building. This is because they feature an almost-flat inflaton direction by construction, yet the other scalar fields couple to the inflaton at the tree level, which allows inflation to naturally end by a second field becoming tachyonic.
@@
We will compute the CMB observables as a function of the model parameters, and study their correlations. We will also present naturalness arguments to identify the preferred region of parameters, and derive the corresponding predictions.
@@
While naturalness would prefer an inflationary scale not too far from the Planck scale, lower-scale scenarios are more interesting for the potential observation of gravitational wave (GW) signals. In hybrid models such as ours, inflation ends because a steep tachyonic “waterfall” direction appears in field space [7], which leads to a transition from the inflationary slow-roll regime to a phase of tachyonic preheating [8, 9]. In this phase, GWs can be produced, and may be within the reach of future detectors, if the inflationary scale is low. In addition, spontaneous symmetry breaking (SSB) in the inflaton-waterfall potential may lead to the production of topological defects, which also source GWs. We will present slightly extended versions of the minimal model, which produce cosmic strings or domain walls after the end of inflation. These scenarios are phenomenologically interesting only if the inflationary scale is sufficiently low, as, in this case, the annihilations of topological defects produce GWs which can be within experimental reach.
@@
Primordial GW sources can potentially explain the recent observation of a stochastic background by the pulsar-timing-array (PTA) collaborations NANOGrav [10], EPTA and InPTA [11], PPTA [12], and CPTA [13]. We will compute the various GW signals in our scenario, study their correlations with the CMB observables, and compare them with present constraints from LIGO-Virgo-KAGRA [14] and NANOGrav [15], as well as with the expected sensitivity of future GW detectors.
@@
Both large-field and small-field models are the subject of several swampland conjectures, which aim to characterize the set of effective field theories that are supposedly incompatible with general principles of quantum gravity; see Ref. [16] for a review. Specifically, models with transplanckian field excursions are concerned by the Distance Conjecture [17]. This conjecture states roughly that, at large distances in field space, an infinite tower of fields becomes light, invalidating the effective field theory. On the other hand, slow-roll potentials with subplanckian field excursions and positive vacuum energy are claimed to be ruled out by various (arguably more speculative) de Sitter conjectures [18, 19, 20]. While these conjectures are certainly interesting, none of them is firmly established. For this work, we, therefore, assume that single-field slow-roll inflation is not per se in the swampland, and that small-field models, in particular, can be viable effective field theories of inflation.
@@ 2 Hybrid inflation with an accidentally light scalar
Here, $T^{a}_{BC}$ are the $\mathrm{SO}(3)$ generators in the ${\bf 5}$-plet representation ($a=1,2,3$ and $A,B,C=1\ldots 5$). We refer the interested reader to Appendix A for a detailed derivation of Eq. (1), and a comprehensive analysis of the vacuum structure of this potential.
The continuous symmetry of $V$ is strictly $G=\mathrm{SO}(3)$: There is no larger custodial-type symmetry present. More precisely, the symmetry of the kinetic terms for $\phi_{A}$ and $\chi_{a}$ is $\mathrm{SO}(8)$, which is explicitly broken to $\mathrm{SO}(5)_{\phi}\times\mathrm{SO}(3)_{\chi}$ by the mass terms and the quartic couplings $\lambda_{\phi,\chi}$ and $\epsilon$. It is non-trivial to show that a single remaining coupling, $\zeta$, breaks explicitly ten generators of this group, reducing the symmetry to the “diagonal” $\mathrm{SO}(3)$, which acts on both $\phi$ and $\chi$. The absence of a larger symmetry implies that, after spontaneous symmetry breaking, there will be at most three Nambu-Goldstone bosons (NGBs); therefore, the appearance of additional, accidentally massless scalars is a non-trivial phenomenon.
@@
At a special point along the flat direction, corresponding to a null eigenvector of one of the $G$ generators, an O$(2)=\mathrm{SO}(2)\rtimes\mathbb{Z}_{2}$ subgroup of $G$ is preserved, where the symbol $\rtimes$ denotes a semidirect product. At this point, there are only two NGBs, while not one but two scalars remain accidentally massless at the tree level; the second one is the scalar which is absorbed by the $\mathrm{SO}(2)$ gauge boson outside the special point. This point turns out to be a minimum of the one-loop effective potential.
For concreteness, consider an explicit parametrization of the tree-level flat direction by writing the isospin-2 field $\phi$ as a $3\times 3$ traceless symmetric matrix $\Phi=\phi_{A}\lambda_{A}$. Here, $\lambda_{A}$ are the the symmetric Gell-Mann matrices $\lambda_{1,3,4,6,8}$. By an $\mathrm{SO}(3)$ transformation, $\Phi$ can be diagonalized, such that its vacuum expectation value (VEV) becomes
@@
$m_{\chi^{1,2}}^{2}=-\tilde{\mu}_{\chi}^{2}+2\,v^{2}\zeta\,\sin^{2}\left(\alpha\pm\frac{\pi}{3}\right)\,,\qquad m_{\chi^{3}}^{2}=-\tilde{\mu}_{\chi}^{2}+2\,\zeta v^{2}\sin^{2}\alpha$ (4)
@@
Let us now consider the effect of one-loop corrections to the potential. The procedure to calculate the one-loop effective potential is detailed in Appendix A. When including one-loop corrections, the vacuum degeneracy is lifted, the physical vacuum is at the point $\alpha=0$, and the two tree-level-massless scalars become massive. The only massless state is the gauge boson of the unbroken $\mathrm{SO}(2)$. However, there remains a one-loop mass hierarchy between the six heavy scalar modes having obtained their masses at the tree level, and the two accidentally light ones with one-loop masses. We emphasize that the latter are calculable and finite as functions of $\mu_{\phi}^{2}$ and $\mu_{\chi}^{2}$, and, therefore, they are quadratically sensitive to quantum corrections at the cutoff scale only at two loops (since contributions to the mass parameters themselves are quadratically divergent at one loop), much as in Little Higgs models (for a review of the latter, see, e.g., Ref. [21]).
Focusing on the one-loop effective potential along the tree-level flat direction, one finds that it is given by  [^1] See Appendix A for the renormalisation prescriptions employed to determine $V_{\rm eff}$, and for the proof that it can be expanded as $\sum_{n}c_{n}\cos(na/f)$.
$V_{\rm eff}(a)=V_{0}-M^{4}\cos\frac{a}{f}+(\text{higher harmonics})\leavevmode\nobreak\ ,$ (6)
@@
$M^{4}=\frac{1}{160\,\pi^{2}}\Bigg{[}9\,g^{4}\,v^{4}+2\,\frac{\tilde{\mu}_{\chi}^{10}}{\zeta^{3}\,v^{6}}\Bigg{(}F\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)-T_{F,5}\left(\frac{\zeta v^{2}}{\tilde{\mu}_{\chi}^{2}}\right)\Bigg{)}\Bigg{]}\leavevmode\nobreak\ ,$ (7)
@@
The shape of the one-loop potential, therefore, closely resembles what one would obtain if $a$ were the pNGB of an approximate global $\mathrm{U}(1)$ symmetry. Identifying $a$ with the inflaton and setting $V_{0}=M^{4}$, one obtains a realization of the famous NI model [4]. However, vanilla NI is excluded by Planck data [2]. Simple modifications of the minimal NI model, such as a non-minimal coupling to gravity [22, 23, 24] or allowing for some exotic phase of reheating [25], can improve the fit, but in the preferred parameter region one finds $f\gtrsim M_{\mathrm{P}}$. [^3] We designate by $M_{\mathrm{P}}$ the reduced Planck mass: $M_{\mathrm{P}}\equiv(8\pi G_{\mathrm{\scriptscriptstyle{N}}})^{-1/2}=2.4\times 10^{18}\,{\rm GeV}$. As already discussed, transplanckian field excursions, apart from being constrained by the Swampland Distance Conjecture, imply that the effects of higher-dimensional operators are not under control (although models have been constructed which address this issue [26, 27]). To overcome these problems, it is desirable to identify a variant of the model where the field excursion is sub-Planckian, i.e., a small-field model of inflation.
In a small-field version of NI, the inflationary potential should be dominated by the constant term $V_{0}$ in the vacuum energy. This implies that slow-roll is possible for $f<M_{\mathrm{P}}$, but it also implies that slow-roll never ends on its own, because the slow-roll parameters never grow to become ${\cal O}(1)$, and that the vacuum energy density $V_{0}$ somehow needs to be converted into radiation eventually. To this end, one may extend the model into one of hybrid inflation [7]: add a second scalar field, the waterfall field, which is heavy during inflation but becomes tachyonic for some value of the inflaton, triggering the end of inflation and the start of preheating.
@@
So far, the role of the field $\chi$ was solely to provide a source for loop corrections to the effective potential which lift the flatness of the inflaton potential. We will now identify $\chi$ with the waterfall field by choosing the model parameters such that the true vacuum of the model is no longer at $\langle\chi\rangle=0$ and $\langle\phi\rangle\neq 0$ but rather at $\langle\phi\rangle=0$ and $\langle\chi\rangle\neq 0$. Our previous discussion of the vacuum structure applies as long as the effective $\chi$ masses, given in Eq. (4), are all positive, regardless of the sign of the mass parameter $\tilde{\mu}_{\chi}^{2}$ in Eq. (5). Even for $\tilde{\mu}_{\chi}^{2}>0$, the effective potential is still given by Eqs. (6) and (7) upon replacing $M^{4}$ by its real part. It is only when a tree-level mass becomes tachyonic that our computation of $V_{\rm eff}$ can no longer be trusted.
@@
In this regime, the inflaton $a$ starts rolling close to $\alpha=\pi/6$ or, equivalently, $a=\pi f$, where all scalars except $a$ have positive tree-level masses. In particular, $\chi$ is stabilized at $\chi=0$. For $\alpha>\alpha_{c}$, the one-loop effective potential Eq. (6) can be trusted, and provides a small slope which the inflaton rolls down. As the inflaton crosses the point $\alpha_{c}$, the tree-level vacuum will no longer be given by $\chi=0$ and Eq. (2), since a tree-level tachyon develops in the $\chi^{3}$ direction. Slow-roll ends as the tachyonic $\chi$ mass grows large, and $\chi$ modes are being excited. Eventually, the potential energy of the inflaton is converted into energy in the $\chi$ sector. In Fig. 1, we provide a sketch of the potential along the $a$ (inflaton) and $\chi^{3}$ (waterfall) directions, for the sake of illustration.
Figure 1: Pictorial representation of the inflationary potential, normalised to the inflationary scale $V_{0}$. Close to $a=\pi f$, where inflation begins, the potential along the $\chi^{3}$ direction (purple lines) has positive and large curvature, stabilizing the inflationary trajectory (red line). At the $a=a_{*}$ the pivot scale crosses the Hubble horizon, $RH=k_{*}$ (with $R$ the scale factor). As the inflaton slow-rolls down its potential, $m_{\chi^{3}}^{2}$ decreases until it becomes 00 at the critical point $a_{c}$, where inflation ends. For $a<a_{c}$, the effective potential cannot be trusted, as the accidentally flat direction is no longer a minimum of the tree-level potential. A value of $a_{*}/f$ ($a_{c}/f$) in the green (blue) region would correspond to a tuning $\gtrsim 10\%$ on the initial (final) point of observable inflation, and, therefore, it is theoretically disfavoured.
Reheating should then proceed via some coupling, such as a Higgs portal coupling, between $\chi$ and the Standard Model; in this work, we will not explore reheating in detail. It is easily checked that, for suitable values of the couplings, the true vacuum of the model is located at
@@
which is the (tree-level) potential energy difference between the inflationary field configuration and the true vacuum. In the following, we will assume that the second term in Eq. (11) is not only subdominant, but even negligible.